OpenStudy (anonymous):

Use the given zero to find all the zeros of the function. (Enter your answers as a comma-separated list. Include the given zero in your answer.) Function: f(x) = 4x^3 + 5x^2 + 36x + 45 Zero: 3i x =

OpenStudy (kinggeorge):

That's \(x=3i\) at the end correct?

OpenStudy (kinggeorge):

If that's one zero, then so must \(x=-3i\). This is a property of complex numbers that's very helpful to memorize. Now you have \[(x+3i)(x-3i)=x^2+9=0\]Now use long division to find the last zero.

OpenStudy (anonymous):

no, 3i is the zero

OpenStudy (anonymous):

Function f(x) = 4x3 + 5x2 + 36x + 45 zero: 3i

OpenStudy (kinggeorge):

Alternatively, you could see a trick, and notice that your final polynomial has a \(4x^3\) term and a \(+45\) term, and \(45/9=5\). So then you could test \(4x+5\) and see if \[(4x+5)(x^2+9)=4x^3+5x^2+36x+45\]

OpenStudy (kinggeorge):

When I write \(x=3i\), it means that \(3i\) is a zero for the polynomial.

OpenStudy (anonymous):

ah ok